Find the domain of f(x)^-1 if f(x)=4x-8

A. x≠ 1/4

B. x≠ -4

C. x≠ 0

D. All real numbers

Answer by jim_thompson5910(35256)

f(x) = 4x-8

y = 4x-8

x = 4y-8

x+8 = 4y

(x+8)/4 = y

y = (x+8)/4

y = (1/4)x+8/4

y = (1/4)x+2

f^-1 (x) = (1/4)x+2

This is a linear equation. So the domain is D. All real numbers